EVALUATE THE DETERMINANT OF THE MATRIX
(6 2) THE ANSWER I GOT WAS -18+-2=-20
(-1 -3) But the answer the teacher gave me was 18+2=-16
and I was wondering how is that not negative and could you please explain to me what exactly is a determinant.

Answer by bucky(2100)

Your teacher made a simple mistake by dropping the minus sign on the 18, and you were also incorrect.
.
To evaluate the two-by-two determinant you multiply along the diagonals ... doing the diagonal
that goes from the left top to the right bottom first, and then the diagonal that goes from
the right top to the left bottom next. Then you subtract the second product from the first
product.
.
For your problem, the first diagonal multiplication is +6 times -3 and the resulting
product is -18. The second diagonal multiplication is +2 times -1 and that product is
-2.
.
Then you subtract the first from the second:
.
-18 - (-2) = -18 + 2 = -16
.
Eventually when you learn to work determinants you will be able to use them to solve
sets of equations. Determinants are a very useful tool. Just hang in there and you will
get around to how to use them pretty soon. Meanwhile, learn the rules that your teacher
is presenting to you and practice them.
.
Good luck and cheers. Hope this helps you ...